Write Maxwell equation for harmonically varying field.


1
Expert's answer
2021-04-19T17:09:03-0400

Harmonically varying field can be written using the Euler's law:

"E=E_0 e^{i \omega t}"

"H=H_0 e^{i \omega t}"

Maxwell's third and forth equations in vacuum:

"\displaystyle \nabla \times E = -\frac{1}{c}\frac{\partial B}{\partial t}"

"\displaystyle \nabla \times B = \frac{1}{c} \frac{\partial E}{\partial t}"


"\displaystyle \frac{\partial B}{\partial t}= i\omega B"

"\displaystyle \frac{\partial E}{\partial t}= i\omega E"


Then Maxwell equations for harmonically varying field looks following

"\displaystyle \nabla \times E = -\frac{1}{c} i \omega B"

"\displaystyle \nabla \times B = \frac{1}{c} i \omega E"